OpenStudy (anonymous):

Use the polynomial to answer the question.21k^4 + 9k^3 - 12k^2

jhonyy9 (jhonyy9):

3k2(7k2 +3k -4)=3k2(7k-4)(k+1) k_1,_2=(-3+/- sqrt9+112)/14=(-3+/-11)/14=-1 and 8/14 =4/7